OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!! Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

You're on your own for Part 1. For Part 2, rewrite the 1st equation as: 2x + y = 11 y = 11 - 2x then substitute for y in the 2nd equation. 3x + 4y = 24 3x + 4(11 - 2x) = 24 3x + 44 - 8x = 24 -5x + 44 = 24 -5x = -20 x = 4 Now put x=4 into either equation to find y... or even easier, remember how we rewrote the 1st equation at the beginning of this problem? y = 11 - 2x y = 11 - 2(4) y = 11 - 8 y = 3 So the solution to the system of equations is (4, 3).